2AlPO4 + 3Mg = 2Al + 1Mg3(PO4)2 Calculate the mass of Sl produced if 0.862 moles og Mg react
Answer #1

Assuming Sl was a typing mistake and what is being asked is the mass of Al(since a and s are adjacent on the keyboard).

Since the coefficient of Mg is 3 and that of Al is 2, the coefficient represent moles, so in the reaction 3 moles of Mg reacts with the other reactant to produce 2 moles of Aluminum.

1 mole of Mg would produce 2/3 moles of Al(unitary method).

0.862 moles of Mg produces (2/3)*0.862 Moles Of Al=0.575 moles Of Al

Molecular Mass oF Al=27g

Total mass produced=no of moles of Al*Molecular Mass

=0.575*27=15.525g of Al